# A Genealogy of Multi-Sensor Foundation Models in Remote Sensing

### v1

@@ Acknowledgements.
- Fuller et al. (2023) Anthony Fuller, Koreen Millard, and James R. Green. 2023. CROMA: Remote Sensing Representations with Contrastive Radar-Optical Masked Autoencoders. doi:10.48550/arXiv.2311.00566 arXiv:2311.00566 [cs].
- Gao et al. (2021) Guangshuai Gao, Qingjie Liu, and Yunhong Wang. 2021. Counting From Sky: A Large-Scale Data Set for Remote Sensing Object Counting and a Benchmark Method. IEEE Transactions on Geoscience and Remote Sensing 59, 5 (May 2021), 3642–3655. doi:10.1109/TGRS.2020.3020555 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
- Gao et al. (2022) Tianyu Gao, Xingcheng Yao, and Danqi Chen. 2022. SimCSE: Simple Contrastive Learning of Sentence Embeddings. doi:10.48550/arXiv.2104.08821 arXiv:2104.08821 [cs].
- Grill et al. (2020) Jean-Bastien Grill, Florian Strub, Florent Altché, Corentin Tallec, Pierre H. Richemond, Elena Buchatskaya, Carl Doersch, Bernardo Avila Pires, Zhaohan Daniel Guo, Mohammad Gheshlaghi Azar, Bilal Piot, Koray Kavukcuoglu, Rémi Munos, and Michal Valko. 2020. Bootstrap your own latent: A new approach to self-supervised Learning. doi:10.48550/arXiv.2006.07733 arXiv:2006.07733 [cs, stat].
- Haldar et al. (2012) Dipanwita Haldar, Anup Das, Shiv Mohan, Om Pal, Ramesh S. Hooda, and Manab Chakraborty. 2012. ASSESSMENT OF L-BAND SAR DATA AT DIFFERENT POLARIZATION COMBINATIONS FOR CROP AND OTHER LANDUSE CLASSIFICATION. Progress In Electromagnetics Research B 36 (2012), 303–321. doi:10.2528/PIERB11071106
- He et al. (2021) Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Dollár, and Ross Girshick. 2021. Masked Autoencoders Are Scalable Vision Learners. doi:10.48550/arXiv.2111.06377 arXiv:2111.06377 [cs].
@@
- He et al. (2015) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2015. Deep Residual Learning for Image Recognition. doi:10.48550/arXiv.1512.03385 arXiv:1512.03385 [cs].
- Helber et al. (2019) Patrick Helber, Benjamin Bischke, Andreas Dengel, and Damian Borth. 2019. EuroSAT: A Novel Dataset and Deep Learning Benchmark for Land Use and Land Cover Classification. doi:10.48550/arXiv.1709.00029 arXiv:1709.00029 version: 2.
- Hoekman and Bouman (1993) D. H. Hoekman and B. A. M. Bouman. 1993. Interpretation of C- and X-band radar images over an agricultural area, the Flevoland test site in the Agriscatt-87 campaign. International Journal of Remote Sensing 14, 8 (May 1993), 1577–1594. doi:10.1080/01431169308953987
- Hou et al. (2024) Yujun Hou, Matias Quintana, Maxim Khomiakov, Winston Yap, Jiani Ouyang, Koichi Ito, Zeyu Wang, Tianhong Zhao, and Filip Biljecki. 2024. Global Streetscapes — A comprehensive dataset of 10 million street-level images across 688 cities for urban science and analytics. ISPRS Journal of Photogrammetry and Remote Sensing 215 (Sept. 2024), 216–238. doi:10.1016/j.isprsjprs.2024.06.023
- Huang et al. (2021) Sha Huang, Lina Tang, Joseph P. Hupy, Yang Wang, and Guofan Shao. 2021. A commentary review on the use of normalized difference vegetation index (NDVI) in the era of popular remote sensing. Journal of Forestry Research 32, 1 (Feb. 2021), 1–6. doi:10.1007/s11676-020-01155-1
- Ignatenko et al. (2020) Vladimir Ignatenko, Pekka Laurila, Andrea Radius, Leszek Lamentowski, Oleg Antropov, and Darren Muff. 2020. ICEYE Microsatellite SAR Constellation Status Update: Evaluation of First Commercial Imaging Modes. In IGARSS 2020 - 2020 IEEE International Geoscience and Remote Sensing Symposium. 3581–3584. doi:10.1109/IGARSS39084.2020.9324531 ISSN: 2153-7003.
- Jain et al. (2022a) Pallavi Jain, Bianca Schoen-Phelan, and Robert Ross. 2022a. Self-Supervised Learning for Invariant Representations from Multi-Spectral and SAR Images. IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing 15 (2022), 7797–7808. doi:10.1109/JSTARS.2022.3204888 arXiv:2205.02049 [cs].
@@
- Jalayer et al. (2025) Sepideh Jalayer, Samira Alkaee Taleghan, Rafael Pires de Lima, Behzad Vahedi, Nick Hughes, Farnoush Banaei-Kashani, and Morteza Karimzadeh. 2025. Enhancing and Interpreting Deep Learning for Sea Ice Charting using the AutoICE Benchmark. Remote Sensing Applications: Society and Environment (2025), 101538.
- Jia et al. (2021) Chao Jia, Yinfei Yang, Ye Xia, Yi-Ting Chen, Zarana Parekh, Hieu Pham, Quoc V. Le, Yunhsuan Sung, Zhen Li, and Tom Duerig. 2021. Scaling Up Visual and Vision-Language Representation Learning With Noisy Text Supervision. doi:10.48550/arXiv.2102.05918 arXiv:2102.05918 [cs].
- Jing et al. (2022) Li Jing, Pascal Vincent, Yann LeCun, and Yuandong Tian. 2022. Understanding Dimensional Collapse in Contrastive Self-supervised Learning. doi:10.48550/arXiv.2110.09348 arXiv:2110.09348.
- Justice et al. (2002) C. O Justice, J. R. G Townshend, E. F Vermote, E Masuoka, R. E Wolfe, N Saleous, D. P Roy, and J. T Morisette. 2002. An overview of MODIS Land data processing and product status. Remote Sensing of Environment 83, 1 (Nov. 2002), 3–15. doi:10.1016/S0034-4257(02)00084-6
- Klemmer et al. (2024) Konstantin Klemmer, Esther Rolf, Caleb Robinson, Lester Mackey, and Marc Rußwurm. 2024. SatCLIP: Global, General-Purpose Location Embeddings with Satellite Imagery. doi:10.48550/arXiv.2311.17179 arXiv:2311.17179 [cs].
- Koroteev (2021) M. V. Koroteev. 2021. BERT: A Review of Applications in Natural Language Processing and Understanding. doi:10.48550/arXiv.2103.11943 arXiv:2103.11943 [cs].
- Krieger et al. (2007) Gerhard Krieger, Alberto Moreira, Hauke Fiedler, Irena Hajnsek, Marian Werner, Marwan Younis, and Manfred Zink. 2007. TanDEM-X: A Satellite Formation for High-Resolution SAR Interferometry. IEEE Transactions on Geoscience and Remote Sensing 45, 11 (Nov. 2007), 3317–3341. doi:10.1109/TGRS.2007.900693
@@
- Li et al. (2019) Liunian Harold Li, Mark Yatskar, Da Yin, Cho-Jui Hsieh, and Kai-Wei Chang. 2019. VisualBERT: A Simple and Performant Baseline for Vision and Language. doi:10.48550/arXiv.1908.03557 arXiv:1908.03557 [cs].
- Li et al. (2024a) Xuyang Li, Danfeng Hong, and Jocelyn Chanussot. 2024a. S2MAE: A Spatial-Spectral Pretraining Foundation Model for Spectral Remote Sensing Data. In 2024 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, Seattle, WA, USA, 27696–27705. doi:10.1109/CVPR52733.2024.02616
- Li et al. (2024b) Zhihao Li, Biao Hou, Siteng Ma, Zitong Wu, Xianpeng Guo, Bo Ren, and Licheng Jiao. 2024b. Masked Angle-Aware Autoencoder for Remote Sensing Images. doi:10.48550/arXiv.2408.01946 arXiv:2408.01946 [cs].
- Liu et al. (2024) Fan Liu, Delong Chen, Zhangqingyun Guan, Xiaocong Zhou, Jiale Zhu, Qiaolin Ye, Liyong Fu, and Jun Zhou. 2024. RemoteCLIP: A Vision Language Foundation Model for Remote Sensing. IEEE Transactions on Geoscience and Remote Sensing 62 (2024), 1–16. doi:10.1109/TGRS.2024.3390838 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
- Liu et al. (2025) Zeping Liu, Fan Zhang, Junfeng Jiao, Ni Lao, and Gengchen Mai. 2025. GAIR: Improving Multimodal Geo-Foundation Model with Geo-Aligned Implicit Representations. doi:10.48550/arXiv.2503.16683 arXiv:2503.16683 [cs].
- Long et al. (2021) Yang Long, Gui-Song Xia, Shengyang Li, Wen Yang, Michael Ying Yang, Xiao Xiang Zhu, Liangpei Zhang, and Deren Li. 2021. On Creating Benchmark Dataset for Aerial Image Interpretation: Reviews, Guidances, and Million-AID. IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing 14 (2021), 4205–4230. doi:10.1109/JSTARS.2021.3070368 Conference Name: IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing.
- Lowe (2004) David G. Lowe. 2004. Distinctive Image Features from Scale-Invariant Keypoints. International Journal of Computer Vision 60, 2 (Nov. 2004), 91–110. doi:10.1023/B:VISI.0000029664.99615.94
@@
- Lu et al. (2025) Siqi Lu, Junlin Guo, James R. Zimmer-Dauphinee, Jordan M. Nieusma, Xiao Wang, Parker VanValkenburgh, Steven A. Wernke, and Yuankai Huo. 2025. Vision Foundation Models in Remote Sensing: A Survey. doi:10.48550/arXiv.2408.03464 arXiv:2408.03464 [cs].
- Mahmood (2014) Ahmed Mahmood. 2014. RADARSAT-1 Background Mission Implementation and Accomplishments. Canadian Journal of Remote Sensing 40, 6 (Nov. 2014), 385–395. doi:10.1080/07038992.2014.999913 Publisher: Canadian Aeronautics and Space Institute _eprint: https://doi.org/10.1080/07038992.2014.999913.
- Mai et al. (2023a) Gengchen Mai, Weiming Huang, Jin Sun, Suhang Song, Deepak Mishra, Ninghao Liu, Song Gao, Tianming Liu, Gao Cong, Yingjie Hu, Chris Cundy, Ziyuan Li, Rui Zhu, and Ni Lao. 2023a. On the Opportunities and Challenges of Foundation Models for Geospatial Artificial Intelligence. doi:10.48550/arXiv.2304.06798 arXiv:2304.06798 [cs].
@@
- Mañas et al. (2021) Oscar Mañas, Alexandre Lacoste, Xavier Giro-i Nieto, David Vazquez, and Pau Rodriguez. 2021. Seasonal Contrast: Unsupervised Pre-Training from Uncurated Remote Sensing Data. doi:10.48550/arXiv.2103.16607 arXiv:2103.16607 [cs].
- Morena et al. (2004) L C Morena, James , K V, , and J. Beck. 2004. An introduction to the RADARSAT-2 mission. Canadian Journal of Remote Sensing 30, 3 (Jan. 2004), 221–234. doi:10.5589/m04-004 Publisher: Canadian Aeronautics and Space Institute _eprint: https://doi.org/10.5589/m04-004.
- Munchak et al. (2020) S. Joseph Munchak, Sarah Ringerud, Ludovic Brucker, Yalei You, Iris de Gelis, and Catherine Prigent. 2020. An Active–Passive Microwave Land Surface Database From GPM. IEEE Transactions on Geoscience and Remote Sensing 58, 9 (Sept. 2020), 6224–6242. doi:10.1109/TGRS.2020.2975477 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Neinavaz et al. (2021) Elnaz Neinavaz, Martin Schlerf, Roshanak Darvishzadeh, Max Gerhards, and Andrew K. Skidmore. 2021. Thermal infrared remote sensing of vegetation: Current status and perspectives. International Journal of Applied Earth Observation and Geoinformation 102 (Oct. 2021), 102415. doi:10.1016/j.jag.2021.102415

### v2

@@ Acknowledgements.
- Fuller et al. (2023) Anthony Fuller, Koreen Millard, and James R. Green. 2023. CROMA: Remote Sensing Representations with Contrastive Radar-Optical Masked Autoencoders. https://doi.org/10.48550/arXiv.2311.00566 arXiv:2311.00566 [cs].
- Gamal et al. (2025) Fawzi Gamal, Tariq Afifa, and Fawzi Gamal. 2025. MoE at Scale: From Modular Design to Deployment in Large-Scale Machine Learning Systems. https://doi.org/10.20944/preprints202504.1313.v1
- Gao et al. (2021) Guangshuai Gao, Qingjie Liu, and Yunhong Wang. 2021. Counting From Sky: A Large-Scale Data Set for Remote Sensing Object Counting and a Benchmark Method. IEEE Transactions on Geoscience and Remote Sensing 59, 5 (May 2021), 3642–3655. https://doi.org/10.1109/TGRS.2020.3020555 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
- Gao et al. (2022) Tianyu Gao, Xingcheng Yao, and Danqi Chen. 2022. SimCSE: Simple Contrastive Learning of Sentence Embeddings. https://doi.org/10.48550/arXiv.2104.08821 arXiv:2104.08821 [cs].
- Girdhar et al. (2023) Rohit Girdhar, Alaaeldin El-Nouby, Zhuang Liu, Mannat Singh, Kalyan Vasudev Alwala, Armand Joulin, and Ishan Misra. 2023. ImageBind: One Embedding Space To Bind Them All. https://doi.org/10.48550/arXiv.2305.05665 arXiv:2305.05665 [cs].
- Grill et al. (2020) Jean-Bastien Grill, Florian Strub, Florent Altché, Corentin Tallec, Pierre H. Richemond, Elena Buchatskaya, Carl Doersch, Bernardo Avila Pires, Zhaohan Daniel Guo, Mohammad Gheshlaghi Azar, Bilal Piot, Koray Kavukcuoglu, Rémi Munos, and Michal Valko. 2020. Bootstrap your own latent: A new approach to self-supervised Learning. https://doi.org/10.48550/arXiv.2006.07733 arXiv:2006.07733 [cs, stat].
- Guo et al. (2024) Xin Guo, Jiangwei Lao, Bo Dang, Yingying Zhang, Lei Yu, Lixiang Ru, Liheng Zhong, Ziyuan Huang, Kang Wu, Dingxiang Hu, Huimei He, Jian Wang, Jingdong Chen, Ming Yang, Yongjun Zhang, and Yansheng Li. 2024. SkySense: A Multi-Modal Remote Sensing Foundation Model Towards Universal Interpretation for Earth Observation Imagery. In 2024 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, Seattle, WA, USA, 27662–27673. https://doi.org/10.1109/CVPR52733.2024.02613
- Haldar et al. (2012) Dipanwita Haldar, Anup Das, Shiv Mohan, Om Pal, Ramesh S. Hooda, and Manab Chakraborty. 2012. ASSESSMENT OF L-BAND SAR DATA AT DIFFERENT POLARIZATION COMBINATIONS FOR CROP AND OTHER LANDUSE CLASSIFICATION. Progress In Electromagnetics Research B 36 (2012), 303–321. https://doi.org/10.2528/PIERB11071106
- He et al. (2021) Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Dollár, and Ross Girshick. 2021. Masked Autoencoders Are Scalable Vision Learners. https://doi.org/10.48550/arXiv.2111.06377 arXiv:2111.06377 [cs].
@@
- He et al. (2015) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2015. Deep Residual Learning for Image Recognition. https://doi.org/10.48550/arXiv.1512.03385 arXiv:1512.03385 [cs].
- Helber et al. (2019) Patrick Helber, Benjamin Bischke, Andreas Dengel, and Damian Borth. 2019. EuroSAT: A Novel Dataset and Deep Learning Benchmark for Land Use and Land Cover Classification. https://doi.org/10.48550/arXiv.1709.00029 arXiv:1709.00029 version: 2.
- Hoekman and Bouman (1993) D. H. Hoekman and B. A. M. Bouman. 1993. Interpretation of C- and X-band radar images over an agricultural area, the Flevoland test site in the Agriscatt-87 campaign. International Journal of Remote Sensing 14, 8 (May 1993), 1577–1594. https://doi.org/10.1080/01431169308953987
- Hong et al. (2024) Danfeng Hong, Bing Zhang, Xuyang Li, Yuxuan Li, Chenyu Li, Jing Yao, Naoto Yokoya, Hao Li, Pedram Ghamisi, Xiuping Jia, Antonio Plaza, Paolo Gamba, Jon Atli Benediktsson, and Jocelyn Chanussot. 2024. SpectralGPT: Spectral Remote Sensing Foundation Model. IEEE Transactions on Pattern Analysis and Machine Intelligence 46, 8 (Aug. 2024), 5227–5244. https://doi.org/10.1109/TPAMI.2024.3362475 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
- Hou et al. (2024) Yujun Hou, Matias Quintana, Maxim Khomiakov, Winston Yap, Jiani Ouyang, Koichi Ito, Zeyu Wang, Tianhong Zhao, and Filip Biljecki. 2024. Global Streetscapes — A comprehensive dataset of 10 million street-level images across 688 cities for urban science and analytics. ISPRS Journal of Photogrammetry and Remote Sensing 215 (Sept. 2024), 216–238. https://doi.org/10.1016/j.isprsjprs.2024.06.023
- Huang et al. (2021) Sha Huang, Lina Tang, Joseph P. Hupy, Yang Wang, and Guofan Shao. 2021. A commentary review on the use of normalized difference vegetation index (NDVI) in the era of popular remote sensing. Journal of Forestry Research 32, 1 (Feb. 2021), 1–6. https://doi.org/10.1007/s11676-020-01155-1
- Ignatenko et al. (2020) Vladimir Ignatenko, Pekka Laurila, Andrea Radius, Leszek Lamentowski, Oleg Antropov, and Darren Muff. 2020. ICEYE Microsatellite SAR Constellation Status Update: Evaluation of First Commercial Imaging Modes. In IGARSS 2020 - 2020 IEEE International Geoscience and Remote Sensing Symposium. 3581–3584. https://doi.org/10.1109/IGARSS39084.2020.9324531 ISSN: 2153-7003.
- Irvin et al. (2023) Jeremy Irvin, Lucas Tao, Joanne Zhou, Yuntao Ma, Langston Nashold, Benjamin Liu, and Andrew Y. Ng. 2023. USat: A Unified Self-Supervised Encoder for Multi-Sensor Satellite Imagery. https://doi.org/10.48550/arXiv.2312.02199 arXiv:2312.02199 [cs].
- Jain et al. (2022a) Pallavi Jain, Bianca Schoen-Phelan, and Robert Ross. 2022a. Self-Supervised Learning for Invariant Representations from Multi-Spectral and SAR Images. IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing 15 (2022), 7797–7808. https://doi.org/10.1109/JSTARS.2022.3204888 arXiv:2205.02049 [cs].
@@
- Jang et al. (2023) Jiho Jang, Seonhoon Kim, Kiyoon Yoo, Chaerin Kong, Jangho Kim, and Nojun Kwak. 2023. Self-Distilled Self-supervised Representation Learning. In 2023 IEEE/CVF Winter Conference on Applications of Computer Vision (WACV). IEEE, Waikoloa, HI, USA, 2828–2838. https://doi.org/10.1109/WACV56688.2023.00285
- Jia et al. (2021) Chao Jia, Yinfei Yang, Ye Xia, Yi-Ting Chen, Zarana Parekh, Hieu Pham, Quoc V. Le, Yunhsuan Sung, Zhen Li, and Tom Duerig. 2021. Scaling Up Visual and Vision-Language Representation Learning With Noisy Text Supervision. https://doi.org/10.48550/arXiv.2102.05918 arXiv:2102.05918 [cs].
- Jing et al. (2022) Li Jing, Pascal Vincent, Yann LeCun, and Yuandong Tian. 2022. Understanding Dimensional Collapse in Contrastive Self-supervised Learning. https://doi.org/10.48550/arXiv.2110.09348 arXiv:2110.09348.
- Justice et al. (2002) C. O Justice, J. R. G Townshend, E. F Vermote, E Masuoka, R. E Wolfe, N Saleous, D. P Roy, and J. T Morisette. 2002. An overview of MODIS Land data processing and product status. Remote Sensing of Environment 83, 1 (Nov. 2002), 3–15. https://doi.org/10.1016/S0034-4257(02)00084-6
- Kaplan et al. (2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B. Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeffrey Wu, and Dario Amodei. 2020. Scaling Laws for Neural Language Models. https://doi.org/10.48550/arXiv.2001.08361 arXiv:2001.08361 [cs].
- Klemmer et al. (2024) Konstantin Klemmer, Esther Rolf, Caleb Robinson, Lester Mackey, and Marc Rußwurm. 2024. SatCLIP: Global, General-Purpose Location Embeddings with Satellite Imagery. https://doi.org/10.48550/arXiv.2311.17179 arXiv:2311.17179 [cs].
- Knight (2008) Philip A. Knight. 2008. The Sinkhorn–Knopp Algorithm: Convergence and Applications. SIAM J. Matrix Anal. Appl. 30, 1 (Jan. 2008), 261–275. https://doi.org/10.1137/060659624 Publisher: Society for Industrial and Applied Mathematics.
- Koroteev (2021) M. V. Koroteev. 2021. BERT: A Review of Applications in Natural Language Processing and Understanding. https://doi.org/10.48550/arXiv.2103.11943 arXiv:2103.11943 [cs].
- Krieger et al. (2007) Gerhard Krieger, Alberto Moreira, Hauke Fiedler, Irena Hajnsek, Marian Werner, Marwan Younis, and Manfred Zink. 2007. TanDEM-X: A Satellite Formation for High-Resolution SAR Interferometry. IEEE Transactions on Geoscience and Remote Sensing 45, 11 (Nov. 2007), 3317–3341. https://doi.org/10.1109/TGRS.2007.900693
@@
- Li et al. (2019) Liunian Harold Li, Mark Yatskar, Da Yin, Cho-Jui Hsieh, and Kai-Wei Chang. 2019. VisualBERT: A Simple and Performant Baseline for Vision and Language. https://doi.org/10.48550/arXiv.1908.03557 arXiv:1908.03557 [cs].
- Li et al. (2024a) Xuyang Li, Danfeng Hong, and Jocelyn Chanussot. 2024a. S2MAE: A Spatial-Spectral Pretraining Foundation Model for Spectral Remote Sensing Data. In 2024 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, Seattle, WA, USA, 27696–27705. https://doi.org/10.1109/CVPR52733.2024.02616
- Li et al. (2024b) Zhihao Li, Biao Hou, Siteng Ma, Zitong Wu, Xianpeng Guo, Bo Ren, and Licheng Jiao. 2024b. Masked Angle-Aware Autoencoder for Remote Sensing Images. https://doi.org/10.48550/arXiv.2408.01946 arXiv:2408.01946 [cs].
- Liu et al. (2024) Fan Liu, Delong Chen, Zhangqingyun Guan, Xiaocong Zhou, Jiale Zhu, Qiaolin Ye, Liyong Fu, and Jun Zhou. 2024. RemoteCLIP: A Vision Language Foundation Model for Remote Sensing. IEEE Transactions on Geoscience and Remote Sensing 62 (2024), 1–16. https://doi.org/10.1109/TGRS.2024.3390838 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
- Liu et al. (2025) Zeping Liu, Fan Zhang, Junfeng Jiao, Ni Lao, and Gengchen Mai. 2025. GAIR: Improving Multimodal Geo-Foundation Model with Geo-Aligned Implicit Representations. https://doi.org/10.48550/arXiv.2503.16683 arXiv:2503.16683 [cs].
- Long et al. (2021) Yang Long, Gui-Song Xia, Shengyang Li, Wen Yang, Michael Ying Yang, Xiao Xiang Zhu, Liangpei Zhang, and Deren Li. 2021. On Creating Benchmark Dataset for Aerial Image Interpretation: Reviews, Guidances, and Million-AID. IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing 14 (2021), 4205–4230. https://doi.org/10.1109/JSTARS.2021.3070368 Conference Name: IEEE Journal of Selected Topics in Applied Earth Observations and Remote Sensing.
- Lowe (2004) David G. Lowe. 2004. Distinctive Image Features from Scale-Invariant Keypoints. International Journal of Computer Vision 60, 2 (Nov. 2004), 91–110. https://doi.org/10.1023/B:VISI.0000029664.99615.94
@@
- Lu et al. (2025) Siqi Lu, Junlin Guo, James R. Zimmer-Dauphinee, Jordan M. Nieusma, Xiao Wang, Parker VanValkenburgh, Steven A. Wernke, and Yuankai Huo. 2025. Vision Foundation Models in Remote Sensing: A Survey. https://doi.org/10.48550/arXiv.2408.03464 arXiv:2408.03464 [cs].
- Mahmood (2014) Ahmed Mahmood. 2014. RADARSAT-1 Background Mission Implementation and Accomplishments. Canadian Journal of Remote Sensing 40, 6 (Nov. 2014), 385–395. https://doi.org/10.1080/07038992.2014.999913 Publisher: Canadian Aeronautics and Space Institute _eprint: https://doi.org/10.1080/07038992.2014.999913.
- Mai et al. (2023a) Gengchen Mai, Weiming Huang, Jin Sun, Suhang Song, Deepak Mishra, Ninghao Liu, Song Gao, Tianming Liu, Gao Cong, Yingjie Hu, Chris Cundy, Ziyuan Li, Rui Zhu, and Ni Lao. 2023a. On the Opportunities and Challenges of Foundation Models for Geospatial Artificial Intelligence. https://doi.org/10.48550/arXiv.2304.06798 arXiv:2304.06798 [cs].
@@
- Mañas et al. (2021) Oscar Mañas, Alexandre Lacoste, Xavier Giro-i Nieto, David Vazquez, and Pau Rodriguez. 2021. Seasonal Contrast: Unsupervised Pre-Training from Uncurated Remote Sensing Data. https://doi.org/10.48550/arXiv.2103.16607 arXiv:2103.16607 [cs].
- Mendieta et al. (2023) Matias Mendieta, Boran Han, Xingjian Shi, Yi Zhu, and Chen Chen. 2023. Towards Geospatial Foundation Models via Continual Pretraining. https://doi.org/10.48550/arXiv.2302.04476 arXiv:2302.04476 [cs].
- Miller (2004) Harvey J. Miller. 2004. Tobler’s First Law and Spatial Analysis. Annals of the Association of American Geographers 94, 2 (June 2004), 284–289. https://doi.org/10.1111/j.1467-8306.2004.09402005.x Publisher: Routledge _eprint: https://doi.org/10.1111/j.1467-8306.2004.09402005.x.
- Morena et al. (2004) L C Morena, James , K V, , and J. Beck. 2004. An introduction to the RADARSAT-2 mission. Canadian Journal of Remote Sensing 30, 3 (Jan. 2004), 221–234. https://doi.org/10.5589/m04-004 Publisher: Canadian Aeronautics and Space Institute _eprint: https://doi.org/10.5589/m04-004.
- Munchak et al. (2020) S. Joseph Munchak, Sarah Ringerud, Ludovic Brucker, Yalei You, Iris de Gelis, and Catherine Prigent. 2020. An Active–Passive Microwave Land Surface Database From GPM. IEEE Transactions on Geoscience and Remote Sensing 58, 9 (Sept. 2020), 6224–6242. https://doi.org/10.1109/TGRS.2020.2975477 Conference Name: IEEE Transactions on Geoscience and Remote Sensing.
@@
- Neinavaz et al. (2021) Elnaz Neinavaz, Martin Schlerf, Roshanak Darvishzadeh, Max Gerhards, and Andrew K. Skidmore. 2021. Thermal infrared remote sensing of vegetation: Current status and perspectives. International Journal of Applied Earth Observation and Geoinformation 102 (Oct. 2021), 102415. https://doi.org/10.1016/j.jag.2021.102415
- Noman et al. (2024) Mubashir Noman, Muzammal Naseer, Hisham Cholakkal, Rao Muhammad Anwar, Salman Khan, and Fahad Shahbaz Khan. 2024. Rethinking Transformers Pre-training for Multi-Spectral Satellite Imagery. https://doi.org/10.48550/arXiv.2403.05419 arXiv:2403.05419 [cs].